# OpenECAD: An Efficient Visual Language Model for Computer-Aided Design

### v2

@@ Abstract
Computer-aided design (CAD) tools are utilized in the manufacturing industry for modeling everything from cups to spacecraft. These programs are complex to use and typically require years of training and experience to master. Structured and well-constrained 2D sketches and 3D constructions are crucial components of CAD modeling. A well-executed CAD model can be seamlessly integrated into the manufacturing process, thereby enhancing production efficiency. Deep generative models of 3D shapes and 3D object reconstruction models have garnered significant research interest. However, most of these models are represented in discrete forms. Moreover, the few models based on CAD operations often have substantial input restrictions. In this work, we fine-tuned pre-trained models to create OpenECAD (0.55B, 0.89B, 2.4B, 3.2B and 4.2B), leveraging the visual, logical, coding, and general capabilities of visual language models. OpenECAD can process images of 3D designs as input and generate highly structured 2D sketches and 3D construction commands. These outputs can be directly used with existing CAD tools’ APIs to generate project files. To train our network, we created a new CAD dataset. This dataset is based on existing public CAD datasets, with adjustments and augmentations to meet the requirements of VLM training.
@@
small language model, visual language model, computer aided design, automated CAD design, 2D-3D shape modeling, geometric deep learning, machine learning
journal: Computer-Aided Design
@@ 1 Introduction
Using 3D shape representation aligns with human intuition. In today’s digital era, computer-aided design (CAD) tools are employed in various industrial fields for 3D shape design, including automotive, aerospace, manufacturing, and architectural design. However, due to drafting conventions and the requirements for shape constraints and edit-ability, 3D shapes are still based on 2D sketches. This approach allows for the meticulous development, relation, and annotation of all design details, mirroring the precision of traditional draftsmen.
@@
A typical 3D part drawing process involves multiple ”Sketch-Extrusion” steps. The relationship between sketch and extrusion is not strictly one-to-one; a single sketch can correspond to multiple extrusions, and one extrusion can be composed of multiple sketches. A sketch consists of multiple closed loops formed by several lines and includes constraints within or between them to ensure the sketch is fully defined. The extrusion operation generates a 3D feature based on the 2D sketch(es). Generally, there is a sequential relationship between ”Sketch-Extrusion” steps, as a sketch’s reference plane may rely on an existing face in the model, and constraints within the Sketch may depend on points or lines already present in the model. As shown in Figure 1, this is a simple process for drawing 3D parts.
Figure 1: A schematic diagram of the process for drawing 3D parts using CAD tools.
@@
There has been extensive research on 3D model generation. However, most of this research focuses on creating computer-discretized forms of 3D shapes, such as 3D point clouds[1, 2, 3, 4, 5], voxelized shapes[6, 7, 8, 9], polygon meshes[10, 11, 12], and levelset fields[13, 14, 15, 16, 17]. This approach neglects the essence of 3D shape design—the drawing process. In practical applications, the generated models often do not fully meet the requirements, and other steps in the production process may necessitate design modifications. Consequently, modifications to the generated models are frequently needed. However, due to the absence of the drawing process, modifying these models can be less efficient than manually modeling from scratch, making these methods challenging to apply in real-world scenarios.
To address this issue, some studies have emerged with the development of neural networks. Since 3D shapes are still based on 2D sketches, these studies can be divided into 2D sketch generation and 3D CAD model generation. SketchGraphs[18] provides a dataset of constrained 2D sketches, CurveGen and TurtleGen[19] generate closed curve sketches, and CADL[20] uses language models to generate sketches. For 3D model generation, SolidGen[21] and ComplexGen[22] generate B-rep models, while Point2Cyl[23], DeepCAD[24], and Free2CAD[25] generate CAD command sequences. It is worth noting that these models often have significant input limitations. For example, they may require existing 3D point clouds, may not accept inputs to limit how to generate, or require isometric hand-drawn sketches. These constraints significantly differ from existing manual modeling workflows, making it challenging to apply them in production processes.
@@
Now, with the development of language models and multimodal language models, these limitations are expected to be overcome. Multimodal language models can flexibly accept various forms of input and understand their meaning, thereby constraining the output. Currently, multimodal language models already have the capability to accept model images and program code and describe their meaning in simple language. For example, GPT-4o can possess simple 3D shape’s generation and has understanding capabilities, and many language models can interpret program code. Therefore, in this work, we leverage the visual, logical, coding, and general capabilities of visual language models to make the following contributions:
@@
    We established OpenECAD dataset. Based on existing public CAD datasets, such as SketchGraphs[18], Fusion 360 Gallery[26], and DeepCAD[24], we convert these datasets into ”image-CAD operation code” pairs, including 2D sketch generation and 3D shape generation. Using currently available multimodal large language models, such as GPT-4o, we generate high-quality annotated datasets. Additionally, we utilize rendering tools to generate various forms of images to enhance the versatility of model inputs.
@@
    We trained OpenECAD model. Based on existing public small language models and multimodal small language models, such as TinyLLaMA[27], OpenELM[28], Gemma[29], Phi-2, Phi-3-mini[30], and TinyLLaVA[31], we train them using the aforementioned training sets and test their performance in CAD design.
@@
    We integrated the code generated by OpenECAD into CAD tools. We design a program that can interface with the APIs of various CAD tools, enabling the generated model code to be directly used in CAD applications.
@@ 3 Method overview
We now present our OpenECAD dataset and model, which includes a new CAD command sequence format designed for language models (see Figure 2). The dataset has been improved based on existing CAD datasets and utilizes large language models to generate high-quality natural language descriptions of CAD models. We fine-tuned pre-trained vision-language models based on small language models, resulting in the OpenECAD models that are compact in size and capable of running locally, specifically designed for CAD applications.
@@ 4 Generation of Datasets
In this chapter, we encompass several key components essential for OpenECAD datasets. We cover the design and generation of datasets (Section 4.1), the definition of the code format for CAD operation sequences (Section 4.2), and the translation of these sequences (Section 4.3). Additionally, we address the generation of natural language descriptions for CAD models (Section 4.4), the production of CAD model view images (Section 4.5), and the overall creation of the CAD dataset (Section 4.6).
4.1 Design and Generation of Datasets
CAD models can be represented in various ways. At the user interaction level with CAD software, 3D shapes are described by a series of CAD operations to create solid forms. For instance, to model a simple table, a user begins by sketching on a 2D plane, drawing a rectangle (a closed curve) and constraining its dimensions. After completing the sketch, the user extrudes it to form the tabletop and constrains its thickness. Next, the user sketches on the existing model, drawing four squares (four closed curves) and constraining their dimensions. After completing this sketch, the user extrudes these shapes to form the table legs, merging them with the previous model and constraining their height. This completes the modeling of a simple table (see Figure 3). This process constitutes a CAD operation sequence, consisting of two Sketch-Extrusion pairs. Clearly, this CAD operation sequence can be described in natural language, as demonstrated above. Additionally, the 2D view of a CAD model is the most intuitive way to express it to humans, akin to the flowchart mentioned above.
Figure 3: A CAD Command Sequence Diagram for Drawing a Table.
In this work, we aim to generate CAD operation sequences using visual language models. CAD operation sequences not only interact with CAD software but also have meanings understandable by humans, allowing modifications and applications in other design processes. To constrain the generation of CAD operation sequences, we use human-understandable inputs such as model images or textual descriptions. Therefore, our dataset includes pairs of target model images as input with CAD operation codes and descriptions as output, pairs of target models as input with CAD operation codes as output, and pairs of target model images and descriptions as input with CAD operation codes as output. These datasets encompass both 3D CAD models and pure 2D sketches to enhance the model’s ability to generate and constrain 2D sketches.
@@ 4.2 Definition of the Code Format for CAD Operation Sequences
Mature CAD tools support a rich set of commands, most of which are simplified calls to collections of basic commands—for example, a rectangle is a collection of four line segments and some constraints. However, in practice, only a small subset of these commands is widely used. Therefore, we currently consider only the most basic and commonly used commands in our dataset (see Table 1). Due to the versatility of language models, extending the dataset later is relatively easy.
@@
add_sketchplane origin_point, x_axis, y_axis
@@
Sketch. Sketch is the core of CAD operations. Extrusion is built upon Sketch. Curves and constraints form closed loops within a Sketch. Multiple closed loops make up the profile of a closed region. In our representation, the definition of Sketch begins with defining the sketch reference plane, similar to most CAD tools. We then maintain a list of loops and start drawing curves and adding constraints. Once a loop is completed, it is stored in the list. After all loops are completed, the corresponding profile is generated, and finally, the sketch is created based on the reference plane and the profile. The sketch reference plane is defined by the three-dimensional coordinates of the origin and the direction of the sketch’s XY axes.
Curves. In practice, we consider three types of widely used curve commands: drawing a line, drawing an arc, and drawing a circle. The line is defined by its starting and ending points, the arc is determined by its starting point, ending point, and midpoint, and the circle is specified by its center and radius.
Extrusion. The extrusion command extrudes a sketch from a 2D plane into a 3D body. The extrusion type can be one-sided, symmetric, or two-sided with respect to the sketch plane. The command also specifies how to merge the newly extruded 3D body with the previously created shape using one of the operations: creating a new body, or joining, cutting, or intersecting with the existing body.
All operation names are chosen to be as meaningful as possible to aid the language model in understanding and using them. With these operation commands, we describe a CAD model as a sequence of executable commands in code. The pseudo code corresponding to a pair of Sketch-Extrusion operations is shown in the Algorithm 1.
Algorithm 1 Pseudo Code Example
$SketchPlane$ = add_sketchplane($origin$, $normal$, $x_{axis}$, $y_{axis}$)
$Loops$ and $Curves$ are lists.
$Line1$ = add_line($start$, $end$)
$Line1$ is appended to $Curves$.
$Loop$ = add_loop($Curves$)
$Loop$ is appended to $Loops$.
$Profile$ = add_profile($Loops$)
$Sketch$ = add_sketch($SketchPlane$, $profile$, $position$, $size$)
$Extrude$ = add_extrude($Sketch$, $operation$, $type$, $extent$)
@@ 4.3 Translation of CAD Operation Sequences
According to the definitions provided in the previous section, we can convert any CAD model covered by the described operations (in Table 1) into a CAD operation sequence as defined by us. However, in practice, different CAD tools or datasets may have slight variations in their definitions of CAD operations (see Figure 4). Therefore, we determine and transform each curve-drawing operation according to its degrees of freedom to match the syntax of the OpenECAD dataset. We have developed scripts to convert other CAD models or datasets into CAD operation sequence codes for training our VLM.
@@ 4.5 Generation of CAD Model View Images
We have introduced how to define and generate CAD model codes and textual descriptions, providing text materials for VLM training. In this section, we will briefly explain how to generate image materials for VLM training, specifically by creating 2D views of CAD models. We first use the PythonOCC[33] tool to write scripts that generate step files (B-Rep format) based on the CAD operation sequence codes. Then, we render the step files using PythonOCC. Finally, we select a direction that forms a random angle with the isometric view direction vector to generate 2D views as image materials.

### v3

# OpenECAD: An Efficient Visual Language Model for Editable 3D-CAD Design

@@ Abstract
Computer-aided design (CAD) tools are utilized in the manufacturing industry for modeling everything from cups to spacecraft. These programs are complex to use and typically require years of training and experience to master. Structured and well-constrained 2D sketches and 3D constructions are crucial components of CAD modeling. A well-executed CAD model can be seamlessly integrated into the manufacturing process, thereby enhancing production efficiency. Deep generative models of 3D shapes and 3D object reconstruction models have garnered significant research interest. However, most of these models produce discrete forms of 3D objects that are not editable. Moreover, the few models based on CAD operations often have substantial input restrictions. In this work, we fine-tuned pre-trained models to create OpenECAD models (0.55B, 0.89B, 2.4B and 3.1B), leveraging the visual, logical, coding, and general capabilities of visual language models. OpenECAD models can process images of 3D designs as input and generate highly structured 2D sketches and 3D construction commands, ensuring that the designs are editable. These outputs can be directly used with existing CAD tools’ APIs to generate project files. To train our network, we created a series of OpenECAD datasets. These datasets are derived from existing public CAD datasets, adjusted and augmented to meet the specific requirements of vision language model (VLM) training. Additionally, we have introduced an approach that utilizes dependency relationships to define and generate sketches, further enriching the content and functionality of the datasets.
@@
small language model; visual language model; computer aided design; geometric deep learning
@@ 1 Introduction
In today’s digital era, computer-aided design (CAD) tools are employed in various industrial fields for 3D shape design, including automotive, aerospace, manufacturing, and architectural design. However, due to drafting conventions and the requirements for shape constraints and edit-ability, 3D shapes are still based on 2D sketches. This approach allows for the meticulous development, relation, and annotation of all design details, mirroring the precision of traditional draftsmen. A typical 3D part drawing process involves multiple “Sketch-Extrusion” steps, where one sketch can correspond to multiple extrusions. A sketch consists of multiple closed loops formed by several lines and includes constraints within or between them to ensure the sketch is fully defined. The extrusion operation generates a 3D feature based on the 2D sketch(es). Generally, there is a sequential relationship between “Sketch-Extrusion” steps, as a sketch’s reference plane may rely on an existing face in the model, and constraints within the sketch may depend on points or lines already present in the model.
@@
There has been extensive research on 3D model generation. Most of this research focuses on creating computer-discretized forms of 3D shapes, such as 3D point clouds[1, 2, 3, 4, 5], voxelized shapes[6, 7, 8, 9], polygon meshes[10, 11, 12], and levelset fields[13, 14, 15, 16, 17]. This approach focuses on the application of precise 3D point cloud models in cyberspace, neglecting the essence of 3D shape design—the drawing process, and therefore is not suitable for the design stage of 3D models. With the development of neural networks, some studies have emerged focusing on generating 3D shapes based on points and surfaces. SolidGen[18] and ComplexGen[19] generate B-rep models, while Point2Cyl[20], DeepCAD[21], and Free2CAD[22] generate CAD command sequences. In practical design workflows, generated 3D models often do not fully meet design requirements and need modifications due to production processes like draft angles and forging. Without the drawing process, these modifications can be less efficient than manually modeling from scratch, making these methods challenging to apply. Additionally, these models often have input limitations, such as requiring existing 3D point clouds or detailed hand-drawn information in isometric views.
@@
Now, with the development of language models and multimodal language models, these limitations are being overcome. Multimodal language models can flexibly accept various forms of input and understand their meaning, thereby constraining the output. Currently, multimodal language models already have the basic ability to accept 2D views or language descriptions of 3D shapes and use CAD operations to draw them. For example, GPT-4o can possess simple 3D shape generation and understanding capabilities, CAD-LLM[23] can manipulate engineering sketches by fine-tuning pre-trained language models to generate parametric CAD.
Figure 1: Example of Extrusion Feature Addition Using an Existing Face as Sketch Reference Plane.
However, current models and datasets often define all points, lines, and surfaces using coordinates within space, such as DeepCAD. This method lacks the dependencies between elements, for instance, a sketch’s reference plane can be defined based on existing points, lines, and surfaces rather than solely relying on the origin and direction vectors. This dependency-based approach to defining 3D shapes, which aligns with human modeling methods, enhances robustness, readability, and editability. The Figure 1 illustrates an example of adding an extrusion feature to an existing geometry, where the new extrusion is created using an existing face as the reference plane for the sketch. Additionally, the vision towers of current multimodal language models, such as OpenAI’s CLIP[24] and Google’s SigLIP [25], tend to focus on macroscopic image analysis, making it difficult to extract fine details, which leads to a significant loss of shape details during generation.
In this work, we leverage the visual, logical, coding, and general capabilities of visual language models to make the following contributions:
@@
    We established OpenECAD datasets. We transform the DeepCAD dataset into pairs of “image-CAD operation code”. Where feasible, we employ reference planes to establish new sketch planes, moving beyond the sole reliance on points and direction vectors. The dimensions within the sketches will be consistent with the absolute space, rather than being scaled as in DeepCAD. These changes enhance the editability of the code. Additionally, we will utilize rendering tools to generate images including single random view, single isometric and transparent view, and a combination of isometric view and three orthographic views, enhancing the versatility of model inputs.
@@
    We trained OpenECAD models using existing small language models and multimodal small language models, including OpenELM[26], Gemma[27] and Phi-2[28]. We employed TinyLLaVA’s [29] training methodology, using OpenAI’s CLIP or Google’s SigLIP as vision tower, and training these models with the LLaVA dataset[30] to imbue them with multimodal capabilities. Subsequently, we fine-tuned both the language models and vision towers of these architectures with OpenECAD datasets and evaluated their performance in CAD design. These models, with the largest containing only 3.1B parameters, demonstrate efficient inference capabilities in single GPU environments.
@@
    We designed a program that can read the generated CAD code and interface with the APIs of PythonOCC, enabling the reconstruction of 3D models from the code and testing the models. Unlike previous models, OpenECAD models require only one or few 2D view(s) to understand 3D shapes, without needing point clouds or Brep files. The generated code is highly editable and aligns with human operations in CAD tools, enabling flexible manual modifications.
@@ 2 Related Work
2.1 Sketch-based 3D Modeling Techniques
In recent years, researchers have made significant progress in the field of automatic conversion from 2D sketches to 3D models. “From sketches to CAM models”[31] proposed a method to directly generate 3D CAM models from 2D sketches by identifying high-level geometric information such as “steps” and “pockets” in 2D line drawings to construct 3D model trees. This approach simplifies traditional CAD/CAM workflows, enabling non-professionals to design and manufacture models. “Extracting datums to reconstruct CSG models”[32] focused on extracting datums from 2D engineering sketches and reconstructing Constructive Solid Geometry (CSG) models through feature extraction, feature tree construction, and datum definition, effectively translating design intent. “Isometric Conversion of Mechanical Sketches”[33] introduced the SFBCM (Sketch Feature-Based Conversion Method), utilizing human perception concepts to convert complex mechanical sketches by progressively detecting and extracting simple features. Lastly, “Method to Automatically Convert Sketches of Mechanical Objects”[34] further expanded the types of sketches that can be processed, defining various sketch faces and features, and proposing a detailed algorithmic process that successfully handled sketches of multiple complex mechanical objects. These studies collectively advanced the development of sketch-based 3D modeling techniques, demonstrating the feasibility of extracting complex 3D features solely from 2D images. Furthermore, they laid a solid theoretical and technical foundation for the subsequent generation of 3D shapes using large neural network models.
@@ 3 Method overview
We now present our OpenECAD datasets and models (see Figure 2), which include a new CAD command sequence format designed for language models. These datasets are converted from the DeepCAD dataset and filtered based on the token count corresponding to the language model, resulting in multiple datasets of varying sizes. Additionally, we utilize large language models to generate high-quality natural language descriptions of OpenECAD codes. Based on multiple small language models of different sizes, we fine-tuned pre-trained vision-language models, ultimately obtaining several OpenECAD models of varying sizes. These models are compact, capable of running locally, and specifically designed for CAD applications. The python code generated by the models can produce CAD project files through corresponding API operators of CAD tools.
@@ 4 Generation of Datasets
In this chapter, we encompass several key components essential for OpenECAD datasets. We cover the design of datasets (Section 4.1), the definition of the code format for CAD operation sequences (Section 4.2), and the translation of these sequences (Section 4.3). Additionally, we address the generation of natural language descriptions for CAD models (Section 4.4), the production of CAD model view images (Section 4.5), and the overall creation of OpenECAD datasets (Section 4.6).
4.1 Design of OpenECAD Datasets
CAD models can be represented in various ways. At the user interaction level with CAD software, 3D shapes are described by a series of CAD operations to create solid forms. For example, if we need to add a cuboid to an existing feature, we first need to specify the reference plane for sketching. Then, we draw a rectangle on the sketch plane. Once the sketch is completed, it forms a closed profile, and the extrude operation is applied to this profile. After the extrusion is completed, the addition of the cuboid is finished. By repeatedly performing this operation, we can create a complex 3D shape (see Figure 1). Clearly, this CAD operation sequence can be described in natural language, as demonstrated above. Additionally, the 2D view of a CAD model is the most intuitive way to express it to humans, akin to the flowchart mentioned above.
In this work, we aim to generate CAD operation sequences using visual language models. CAD operation sequences not only interact with CAD software but also have meanings understandable by humans, allowing modifications and applications in other design processes. To constrain the generation of CAD operation sequences, we use human-understandable inputs such as model images or textual descriptions.
@@ 4.2 Definition of the Code Format for CAD Operation Sequences
add_sketchplane origin, normal, x_axis
add_sketchplane_ref extrude, origin, type, (optional values)
@@
Mature CAD tools support a rich set of commands, most of which are simplified calls to collections of basic commands—for example, a rectangle is a collection of four line segments and some constraints. However, in practice, only a small subset of these commands is widely used. Therefore, we currently consider only the most basic and commonly used commands in our dataset (see Table 1).
Curve operations include add line, arc, or circle, all of which use the most intuitive variables for definition (as shown in the definition method on the left side of Figure 3).
Sketch operations include defining the reference plane directly using the origin, direction vector, and x-axis direction, or using existing features to define the reference plane. When using existing features, the inputs vary depending on the type, including: using the same reference plane as the reference extrusion, using the top face or the side face obtained from the extrusion. When using the side face, the corresponding line must be specified. These methods can determine whether to reverse the direction vector or rotate the x-axis. Sketch operations also include auxiliary functions such as obtaining the profile and the sketch.
Extrusion operation requires inputs such as the sketch, operation type, extrusion type, and extrusion length. The operation types include simple extrusion, cut, or join with existing features, and the extrusion types include unidirectional or bidirectional extrusion.
Due to the versatility of language models, extending the dataset later is relatively easy. All operation names are chosen to be as meaningful as possible to aid the language model in understanding and using them. With these operation commands, we describe a CAD model as a sequence of executable commands in code. An example of OpenECAD code is shown in Algorithm 1.
Algorithm 1 OpenECAD Code Example
1:  $SketchPlane0$ = add_sketchplane($origin$, $normal$, $x_{axis}$)
2:  $Loops0$, $Curves0\_0$ = [], []
3:  $Line0\_0\_0$ = add_line($start$, $end$)
4:  …
5:  $Loop0\_0$ = add_loop($Curves0\_0$)
6:  $Profile0$ = add_profile($Loops0\_0$)
7:  $Sketch0$ = add_sketch($SketchPlane0$, $Profile0$, $position$, $size$)
8:  $Extrude0$ = add_extrude($Sketch0$, $operation$, $type$, $extent$)
9:  $SketchPlane1$ = add_sketchplane_ref($Extrude0$, $origin$, $type$= “line”, $line$= $Line0\_0\_1$)
10:  …
@@ 4.3 Translation of CAD Operation Sequences
According to the definitions provided in the previous section, we can convert any CAD model covered by the described operations (in Table 1) into a CAD operation sequence as defined by us. However, in practice, different CAD tools or datasets may have slight variations in their definitions of CAD operations.
For curve operations, we determine and transform each curve-drawing operation according to its degrees of freedom to match the syntax of the OpenECAD dataset. For the DeepCAD dataset, we do not need to convert their definition methods; we only need to convert them into OpenECAD codes to train our vision language model (VLM). However, for other datasets such as AutoDesk Fusion, we need to design scripts to calculate the corresponding relationships based on geometry (see Figure 3).
Algorithm 2 Find reference plane
1:  $t$ is target plane’s normal vector
2:  for all existed extrudes: $E$ do
3:     $n$ is $E$’s normal vector
4:     if $n$ is parallel with $t$ then
5:        if two plane’s distance is 0 or extent_size of $E$ then
6:           reference plane found.
7:           return
8:        else
9:           $E$ can’t contain reference plane.
10:        end if
11:     else if $n$ is perpendicular with $t$ then
12:        for all lines in sketch of $E$: $l$ do
13:           if $l$ on the target plane then
14:              reference plane found.
15:              return
16:           end if
17:        end for
18:     else
19:        $E$ doesn’t contain reference plane.
20:     end if
21:  end for
For sketch operations, the primary focus is on defining the sketch reference plane, and the functions that assist in constructing the sketch do not require special conversion operations. For the DeepCAD dataset, for planes that can only be defined directly using the origin and vectors, we do not need to convert their definition methods, as the direction vector and x-axis can be easily calculated. We only need to convert them into OpenECAD codes. However, for reference planes that can be defined using existing features, we need to design an algorithm to find the appropriate plane. We will traverse all previously defined sketch planes and the planes obtained from extrusions (limited to flat surfaces, excluding curved surfaces) to find a suitable reference plane. This algorithm is shown in Algorithm 2.
For extrusion operations, no special conversion is required.
@@ 4.5 Generation of CAD Model View Images
We have introduced how to define and generate CAD model codes and textual descriptions, providing text materials for VLM training. In this section, we will briefly explain how to generate image materials for VLM training, specifically by creating 2D views of CAD models.
We first use the PythonOCC tool to write scripts that generate step files (B-Rep format) based on the CAD operation sequence codes. Then, we generate three types of 3D shape’s images: Default View, Transparent View, and Orthographic Views. The Default View is a 2D view of the 3D shape generated by selecting a direction that forms a random angle with the isometric view direction vector. The Transparent View makes the 3D shape transparent to reveal features that are not visible from a certain direction. The Orthographic Views consist of a single image composed of the front view, top view, left view, and isometric view. We use PythonOCC to render these step files and generate the corresponding images.